Identify the slope and y-intercept in the equation −3x + 8y = −24.

Answer by checkley77(12844)

Y=mX+b where m=slope & b=Y intercept.
−3x + 8y = −24
8Y=3X-24
Y=3X/8-24/8
Y=3X/8-3 m=(3/8) & b=(-3) is the Y intercept.